### v1

@@ 2 $\pi^{2}$-NN Networks
To leverage interconnects for computing, we study $\pi^{2}$-NN, a type of neural network, consisting of neurons that are arranged in N hidden layers. The events of neurons in layer n1n-1 are sent to neurons in layer $n$. The layers are either fully connected (i.e., each neuron receives input from all neurons in the previous layer) or convolutional (i.e., connections are limited to be local and share weights). In the following equations, an upper index refers to the layer number. The $D$-dimensional real-valued inputs $X$, such as pixel intensities, are encoded into differential event spike times as $T^{0}_{+}=|A+X|_{+},T^{0}_{-}=|A-X|_{+}$ and fed as an input to $\pi^{2}-NN$. The synaptic weights (W) are encoded as differential synaptic delays represented as $W_{+}=|B+W|_{+},W_{-}=|B-W|_{+}$ for all layers. Here, A and B are arbitrary constants, and $|.|=max(0,.)$ ensures that the information represented in time is non-negative. Given the differential spike times and synaptic delays of the neurons in the previous layer (n1)(n-1), the differential input spike times for the next layer $(n)$ are computed using the $K^{th}$ neuron model dynamics (Eq. 25).
$T^{n}_{+},T^{n}_{-}=K^{th}(T^{n-1}_{+},T^{n-1}_{-},W^{n}_{+},W^{n}_{-})$ (25)
2.0.1 An alternate representation of the $\pi^{2}_{K}$ neuron model
The differential formulation of the $K^{\text{th}}$ neuron model, inspired from [23], is as follows: Let $T_{+}^{n-1},\ T_{-}^{n-1},\ W_{+}^{n},\ W_{-}^{n}$ be the differential input spike times and synaptic delays to the $K^{\text{th}}$ neuron at layer $n$.
@@ 2.0.1 An alternate representation of the $\pi^{2}_{K}$ neuron model
    $H1^{n}$ $=\text{sort}[T_{+}^{n-1}+W_{+}^{n},\;T_{-}^{n-1}+W_{-}^{n}]\quad$ $H2^{n}=\text{sort}[T_{+}^{n-1}+W_{-}^{n},\;T_{-}^{n-1}+W_{+}^{n}]$ (26)
@@
    $T_{+}^{n}$ $=M+(H_{-}^{n}-H_{+}^{n})\quad$ $T_{-}^{n}=M+(H_{+}^{n}-H_{-}^{n})$ (27)
Here, K decides the number of inputs that contribute. $K\leq D^{n-1}$, if the output dimension of the n1n-1 layer is is $D^{n-1}$-dimensional. M is a positive constant chosen to maintain the causality between the input and output spike times (i.e. $T_{+}^{n}\ \&\ T_{-}^{n}$ are always greater than the first $K$ inputs that contributed to them). $\alpha$ is a positive scaling constant. These differential outputs are sent as inputs to the next layer. The function can be summarized by Eq. 25. The fundamental operations of the $\pi^{2}$-NN type of neural networks are delaying (each input event is delayed based on its corresponding synaptic weight), time-based sorting (only the first $K$ earliest arriving events are selected for computation), and event dropping ( The remaining (DK)(D-K) later events are discarded). The neurons in the $\pi^{2}$ network accumulate and compute the mean spike time of the first K events that arrive at it. This establishes a neural network design that leverages the inherent properties of interconnects to perform computation.
@@ 2.1 $\pi^{2}$-NN can learn
Figure 10: A) The spiking activity of a 2 x 10 x 2 network trained on the XOR dataset evaluated across input classes and values of K. The coding scheme is spatiotemporal: information is carried by the subset of neurons that activate in each layer and, within that subset, by their temporal order of spiking. Because synaptic delays allow a single neuron to be delayed by multiple offsets, a given unit can participate in multiple contributing groups that drive the next layer. Selectivity, therefore, arises from spatial sparsity (which units participate) and temporal ranking (when they arrive). This joint code is inherently combinatorial, as the number of distinct patterns grows with both the set of active neurons and their permutations in time. A,B represent the spiking pattern for classes 0 and 1 when the K is set to be 2 and 3 at the hidden and output layer, respectively. C,D represent the spiking activity obtained when K is 1 in both layers. All the neurons are represented in a differential manner to depict their differential spiking time instants as presented in Eq. 27. The 2 input neurons $I_{0},I_{1}$ are represented differentially as $I_{0}^{+},I_{1}^{+},I_{0}^{-},I_{1}^{-}$. Similarly, the hidden nodes $H_{0}^{+},H_{1}^{+},...,H_{9}^{+},H_{1}^{-},H_{2}^{-},...,H_{9}^{-}$ are represented using indices 5-24, and the two output nodes for classes 0 and 1 $O_{0}^{+},O_{1}^{+},O_{0}^{-},O_{1}^{-}$ are indexed from 25-29. With K=[2,3] the network achieves a test accuracy of 99.5% and with K set as [1,1] the network achieves a test accuracy of 100%.
@@
Figure 11: The spatio-temporal patterns emerging from the layers of the trained $\pi^{2}$ network are illustrated as a raster plot. The blue and green dots represent the differential event generation times of the nodes in a layer. The activity across a population of neurons (across layers) over time for the classes 1 (A,B) and 4 (C,D) iv) Represents the original image from the MNIST dataset.
@@
Learning in $\pi^{2}$-NN requires the ability to relate the differential output spike times with both the differential input spike times and synaptic delays. In the $K^{th}$ neuron model, the contributions to $H_{+},H_{-}$ come only from the $K$ earliest events. Consequently, the gradient of the output spike times of layer $n$ with respect to the differential weights can be computed as:
$\frac{\partial T_{\text{out}+}^{n}}{\partial W_{i+}^{n}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i+}^{n}\in H1^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i+}^{n}\in H2^{n}[1{:}K])$ (28)
$\frac{\partial T_{\text{out}-}^{n}}{\partial W_{i+}^{n}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i+}^{n}\in H2^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i-}^{n}\in H1^{n}[1{:}K])$ (29)
$\frac{\partial T_{\text{out}+}^{n}}{\partial W_{i-}^{n}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i-}^{n}\in H1^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i-}^{n}\in H2^{n}[1{:}K])$ (30)
$\frac{\partial T_{\text{out}-}^{n}}{\partial W_{i-}^{n}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i-}^{n}\in H2^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i-}^{n}\in H1^{n}[1{:}K])$ (31)
@@
$\frac{\partial T_{\text{out}+}^{n}}{\partial T_{i+}^{n-1}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i+}^{n}\in H1^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i-}^{n}\in H2^{n}[1{:}K])$ (32)
$\frac{\partial T_{\text{out}-}^{n}}{\partial T_{i-}^{n-1}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i-}^{n}\in H1^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i+}^{n}\in H2^{n}[1{:}K])$ (33)
Here, $\mathbbm{1}(J)$ is the indicator function on condition J, which evaluates to 1 if the condition J is met and to 0 otherwise. K and $\alpha$ are hyperparameters for the $K^{th}$ neuron model. Using the exact spike-based derivatives provided by the $K^{\text{th}}$ neuron model, the chain rule can be applied to compute gradients of the loss with respect to the differential weight parameters across a multi-layered $\pi^{2}$-NN.
We can now apply the findings above to study learning in a layered network. The output of the network is defined by the identity of the label neuron that has the maximum difference between its differential output spiking times. During training, the objective is to maximize the temporal separation between the differential output of the correct class neuron than that of all other neurons in the output layer. To facilitate supervised learning, we employ the cross-entropy loss function, applied to the difference of the differential outputs of the neurons, allowing the use of backpropagation to update the delay-encoded synaptic weights throughout all layers of the $\pi^{2}$-NN. Gradient descent on the loss function equation can be easily performed by repeated application of the chain rule, using the exact derivatives in equations 28-33. Using the derived gradients, we train a 2 x 10 x 2 - $\pi^{2}$ network on the XOR dataset, and represent its spiking activity in Fig.10. We choose a small toy problem to better understand and illustrate the spatiotemporal spiking dynamics of $\pi^{2}$ networks. We generate 1,0001{,}000 two-dimensional samples by drawing each coordinate uniformly from [1,1][-1,1] and assign class labels according to the XOR rule on the coordinate signs (points in opposite-sign quadrants are labeled class 1; same-sign quadrants class 0). The dataset is randomly permuted and split $80/20$ into training and test sets with balanced classes.
@@ 3.1 Vision for large-scale systems
Figure 12: A) A subpool of h neurons is fully connected to another subpool of h neurons. B) A port is shared by m neurons. C) The proposed $\pi^{2}$-NN implementation using an Ethernet switch. The Ethernet switch can support $n=mH_{out}$ neurons in total.
We strategize and analyze an implementation for a processing system consisting of n neurons, communicating through an Ethernet switch (with H ports each supporting a bandwidth B), as shown in Fig. 12 C. $m$ neurons share a port. There are sub-pools (among these m neurons), each consisting of h neurons. Thus, a port is shared by m/h sub-pools. We propose an architecture where a sub-pool of h neurons is fully connected to another sub-pool of h neurons. Each sub-pool-pool connection consists of $h^{2}$ synapses. For each of these synaptic connections, $\log_{2}(w)$ bits, representing the synaptic delays, have to be stored. The number of sub-pools in the core is $n/h$. The neuron’s address consists of the port address it corresponds to ($\log_{2}(n/m)\ \text{ bits}$), the pool it belongs to ($\log_{2}(m/h)\ \text{ bits}$), distinct neuron address in that pool ($\log_{2}(h)\ \text{bits}$) - which amounts to a total of $\log_{2}(n)\ \text{ bits}$. Features of this system include:
@@
  • The TX address is decoded, the routing table is accessed, and events with appropriate destination addresses (Rx address) are generated. Similar to the previous implementation, a PCP tag (p bits) is added to the events representing the synaptic delay value. The events are directed to their corresponding destination ports by the switching logic.
  • Every output port is associated with $m$ set of shaped queues (Fig. 2f), one for each neuron that shares that port. A set of shaped queues consists of $2^{p}$ queues each, one for each distinct PCP value. Events associated with the same PCP tags for a neuron are dropped in the same queue. So, the events will inherently be arranged in the ascending order of their arrival times. Thus, the queue selection is based on the Rx neuron address and the PCP tag.
  • The ATS shaper controls each shaped queue. Based on the synaptic delay (PCP tag), a TET tag is assigned to the events to schedule its transmission to the shared queue of the neuron by the ATS shaper. When the appropriate TET value of an event is reached, it is released to the shared queue of that Rx neuron.
  • A modified CBS shaper controls the shared queue of the neuron. As events arrive, the credits are accumulated. When the CBS-controlled gate opens for transmission, an event encoded with the Rx neuron address, consisting of $log_{2}(n)$ bits is forwarded to the input ports of the switch. The remaining events are dropped (TTFS encoding). This event now becomes the next Tx event processed by the switch.
@@
  • For the power analysis, we assume the weights are uniformly distributed (a similar assumption was taken in [73]).
Bandwidth requirements at the input/output port: On average, the number of neurons (m) that can share an port or channel that can support a bandwidth of B bits/s can be computed by solving the following inequality.
$B\geq F_{spk}mlog_{2}(n)$ (34)
Memory requirements: Say we have a look-up table in the switch storing the connectivity information of every pool-pool pair and their corresponding synaptic delays. Here we assume that every input port is associated with a memory block storing the connectivity information of the m neurons that share that port. Once a TX event arrives at an input port, its pool number is extracted ($log_{2}(m/h)\ bits$), and the LUT is accessed to retrieve the destination port ($log_{2}(n/m)\ bits$) and the destination pool address ($log_{2}(m/h)\ bits$) it is associated with. The TX event is fully connected to every neuron in the destination pool. For every pool-pool connection, we have to store $h^{2}log_{2}(w)$ bits. Thus, for one pool-pool connection, the LUT stores $h^{2}log_{2}(w)+log_{2}(nm/h^{2})\ bits$. Every port is associated with m/h pools and there are $H_{in}$ ports. Thus, the total memory required to store the LUT for all the input ports will be:
$M\geq nhlog_{2}(w)+(n/h)log_{2}(mn/h^{2})\ bits$ (35)
@@
$Q_{nsize}=nKlog_{2}(n)F_{spk}T_{out}\ bits(time\ units/s)$ (36)
@@
$QN_{size}=T_{avg}KF_{spk}(log_{2}(n)+n_{t})\ bits(time\ units/s)$ (38)
Here, $n_{t}$ is the number of bits requited to store the TET (transmission elibility time). If more than $KF_{spk}\ events/s$ occur, then they will be automatically dropped. If the synaptic delay values are uniformly distributed among the $2^{p}$ queues of a neuron, then the queue size can be further reduced by $2^{p}$ times. Under this assumption, the total queue memory required in bytes across all the output ports is:
$Q_{size}=T_{avg}nKF_{spk}(log_{2}(n)+n_{t})/(8*2^{p})\ B(time\ units/s)$ (39)
Throughput: Neuromorphic systems are usually compared in terms of synaptic operations per second (SOPS). The SOPS for this system is related to its network activity and can be computed as a product of the average firing rate of the neurons and the average number of active synapses [74, 21]. The throughput of the $\pi^{2}$ system can be computed based on its three fundamental operations: number of effective delay operations performed, number of packets transmitted by the interconnect system (neuron activity), number of packet drops at the queues of the output ports and at the neuron.
@@
It is important to note that, these operations are primarily implemented using the interconnect system (ethernet switches, queues).
@@
Power analysis: The switches, interconnect transmission power and memory access dominate the system’s network power consumption. If the interconnect modules to the switch consume aJ/bita\ J/bit then the interconnect power can be calculated as:
$P_{i}=aBK\ Watts$ (44)
@@
If it takes bJ/bitb\ J/bit to access the memory related to the routing table, and queues at the output port, then the power consumption to access Vbits/sV\ bits/s is
$P_{m}=bV\ Watts$ (45)
@@
The total power of the system consisting of a switch, memory and interconnects (excluding the neurons) can be computed as:
$P_{total}=P_{s}+P_{i}+P_{m}\ Watts$ (46)

### v3

@@ 2 $\pi^{2}$-NN Networks
To leverage interconnects for computing, we study $\pi^{2}$-NN, a type of neural network, consisting of neurons that are arranged in N hidden layers. The events of neurons in layer n1n-1 are sent to neurons in layer $n$. The layers are either fully connected (i.e., each neuron receives input from all neurons in the previous layer) or convolutional (i.e., connections are limited to be local and share weights). In the following equations, an upper index refers to the layer number. The $D$-dimensional real-valued inputs $X$, such as pixel intensities, are encoded into differential event spike times as $T^{0}_{+}=|A+X|_{+},T^{0}_{-}=|A-X|_{+}$ and fed as an input to $\pi^{2}-NN$. The synaptic weights (W) are encoded as differential synaptic delays represented as $W_{+}=|B+W|_{+},W_{-}=|B-W|_{+}$ for all layers. Here, A and B are arbitrary constants, and $|.|=max(0,.)$ ensures that the information represented in time is non-negative. Given the differential spike times and synaptic delays of the neurons in the previous layer (n1)(n-1), the differential input spike times for the next layer $(n)$ are computed using the $\pi^{2}_{K}$ neuron model dynamics (Eq. S6).
$T^{n}_{+},T^{n}_{-}=\pi^{2}_{K}(T^{n-1}_{+},T^{n-1}_{-},W^{n}_{+},W^{n}_{-})$ (S6)
2.1 Algorithmic Realization of the $\pi^{2}_{K}$ Neuron in Software
The differential formulation realization of the $\pi^{2}_{K}$ neuron model, inspired from [temp], is as follows: Let $T_{+}^{n-1},\ T_{-}^{n-1},\ W_{+}^{n},\ W_{-}^{n}$ be the differential input spike times and synaptic delays to the $\pi^{2}_{K}$ neuron at layer $n$.
@@ 2.1 Algorithmic Realization of the $\pi^{2}_{K}$ Neuron in Software
    $H1^{n}$ $=\text{sort}[T_{+}^{n-1}+W_{+}^{n},\;T_{-}^{n-1}+W_{-}^{n}]\quad$ $H2^{n}=\text{sort}[T_{+}^{n-1}+W_{-}^{n},\;T_{-}^{n-1}+W_{+}^{n}]$ (S7)
@@
    $T_{+}^{n}$ $=M+max(0,(H_{-}^{n}-H_{+}^{n}))\quad$ $T_{-}^{n}=M-max(0,(H_{-}^{n}-H_{+}^{n}))$ (S8)
K decides the number of inputs that contribute, and $K\leq D^{n-1}$, if the output dimension of the n1n-1 layer is $D^{n-1}$-dimensional. M is a positive constant chosen to maintain the causality between the input and output spike times (i.e., $T_{+}^{n}\ \&\ T_{-}^{n}$ are always greater than the first $K$ inputs that contributed to them). $\alpha$ is a positive scaling constant. These differential outputs are sent as inputs to the next layer. The function can be summarized by Eq. S6. The fundamental operations of the $\pi^{2}$-NN type of neural networks are delaying (each input event is delayed based on its corresponding synaptic weight), time-based sorting (only the first $K$ earliest arriving events are selected for computation), and event dropping ( The remaining (DK)(D-K) later events are discarded). The neurons in the $\pi^{2}$ network accumulate and compute the mean spike time of the first K events that arrive at it. This establishes a neural network design that leverages the inherent properties of interconnects to perform computation.
@@ 2.2 $\pi^{2}$-NN can learn
Figure S2: A) The spiking activity of a 2 x 10 x 2 network trained on the XOR dataset was evaluated across input classes and values of K. The coding scheme is spatiotemporal: information is carried by the subset of neurons that activate in each layer and, within that subset, by their temporal order of spiking. Because synaptic delays allow a single neuron to be delayed by multiple offsets, a given unit can participate in multiple contributing groups that drive the next layer. Selectivity, therefore, arises from spatial sparsity (which units participate) and temporal ranking (when they arrive). This joint code is inherently combinatorial, as the number of distinct patterns grows with both the set of active neurons and their permutations in time. A,B represent the spiking pattern for classes 0 and 1 when the K is set to be 2 and 3 at the hidden and output layer, respectively. C,D represent the spiking activity obtained when K is 1 in both layers. All the neurons are represented in a differential manner to depict their differential spiking time instants as presented in Eq. S8. The 2 input neurons $I_{0},I_{1}$ are represented differentially as $I_{0}^{+},I_{1}^{+},I_{0}^{-},I_{1}^{-}$. Similarly, the hidden nodes $H_{0}^{+},H_{1}^{+},...,H_{9}^{+},H_{1}^{-},H_{2}^{-},...,H_{9}^{-}$ are represented using indices 5-24, and the two output nodes for classes 0 and 1 $O_{0}^{+},O_{1}^{+},O_{0}^{-},O_{1}^{-}$ are indexed from 25-29. With K=[2,3] the network achieves a test accuracy of 99.5% and with K set as [1,1] the network achieves a test accuracy of 100%.
@@
Figure S3: The spatio-temporal patterns emerging from the layers of the trained $\pi^{2}$ network are illustrated as a raster plot. The blue and green dots represent the differential event generation times of the nodes in a layer. The activity across a population of neurons (across layers) over time for the classes 1 (A,B) and 4 (C,D) iv) represents the original image from the MNIST dataset.
@@
Learning in $\pi^{2}$-NN requires the ability to relate the differential output spike times with both the differential input spike times and synaptic delays. In the $\pi^{2}_{K}$ neuron model, the contributions to $H_{+},H_{-}$ come only from the $K$ earliest events. Consequently, the gradient of the output spike times of layer $n$ with respect to the differential weights can be computed as:
$\frac{\partial T_{+}^{n}}{\partial W_{i+}^{n}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i+}^{n}\in H1^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i+}^{n}\in H2^{n}[1{:}K])$ (S9)
$\frac{\partial T_{-}^{n}}{\partial W_{i+}^{n}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i+}^{n}\in H2^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i-}^{n}\in H1^{n}[1{:}K])$ (S10)
$\frac{\partial T_{+}^{n}}{\partial W_{i-}^{n}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i-}^{n}\in H1^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i-}^{n}\in H2^{n}[1{:}K])$ (S11)
$\frac{\partial T_{-}^{n}}{\partial W_{i-}^{n}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i-}^{n}\in H2^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i-}^{n}\in H1^{n}[1{:}K])$ (S12)
@@
$\frac{\partial T_{+}^{n}}{\partial T_{i+}^{n-1}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i+}^{n}\in H1^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i+}^{n-1}+W_{i-}^{n}\in H2^{n}[1{:}K])$ (S13)
$\frac{\partial T_{-}^{n}}{\partial T_{i-}^{n-1}}$ $=-\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i-}^{n}\in H1^{n}[1{:}K])+\frac{\alpha}{K}\mathbbm{1}(T_{i-}^{n-1}+W_{i+}^{n}\in H2^{n}[1{:}K])$ (S14)
Here, $\mathbbm{1}(J)$ is the indicator function on condition J, which evaluates to 1 if the condition J is met and to 0 otherwise. K and $\alpha$ are hyperparameters for the $\pi^{2}_{K}$ neuron model. Using the exact spike-based derivatives provided by the $\pi^{2}_{K}$ neuron model, the chain rule can be applied to compute gradients of the loss with respect to the differential weight parameters across a multi-layered $\pi^{2}$-NN.
We can now apply the findings above to study learning in a layered network. The output of the network is defined by the identity of the label neuron that has the maximum difference between its differential output spiking times. During training, the objective is to maximize the temporal separation between the differential output of the correct class neuron than that of all other neurons in the output layer. To facilitate supervised learning, we employ the cross-entropy loss function, applied to the difference of the differential outputs of the neurons, allowing the use of backpropagation to update the delay-encoded synaptic weights throughout all layers of the $\pi^{2}$-NN. Gradient descent on the loss function equation can be easily performed by repeated application of the chain rule, using the exact derivatives in equations S9-S14. Using the derived gradients, we train a 2 x 10 x 2 - $\pi^{2}$ network on the XOR dataset, and represent its spiking activity in Fig.S2. We choose a small toy problem to better understand and illustrate the spatiotemporal spiking dynamics of $\pi^{2}$ networks. We generate 1,0001{,}000 two-dimensional samples by drawing each coordinate uniformly from [1,1][-1,1] and assigning class labels according to the XOR rule on the coordinate signs (points in opposite-sign quadrants are labeled class 1; same-sign quadrants class 0). The dataset is randomly permuted and split $80/20$ into training and test sets with balanced classes.
@@ 3 Processing using Ethernet
3.1 $\pi^{2}$-CBS and $\pi^{2}$-ATS Protocol
The required modifications to the CBS protocol (IEEE 802.1Qav) are explicitly detailed in Fig. S4 and Table S1, and the corresponding adaptations to the ATS protocol (IEEE 802.1Qcr) are described in Algorithm 1. When the $\pi^{2}$-CBS and $\pi^{2}$-ATS algorithms are used as described, we observe a 2% accuracy drop relative to SOTA on a LeNet-5 network trained on FMNIST. Introducing the additional scheduling mechanism described in Eq. 9,10 of the manuscript, followed by differential encoding, restores SOTA performance. The primitives required for this scheduling mechanism—ReLU, addition, and subtraction—correspond to well-established scheduling primitives already present in Ethernet switches[vc]. Future work will explore how existing scheduling protocols can be leveraged to improve accuracy and how $\pi^{2}$ networks can be trained to achieve high accuracy without relying on this additional scheduling logic. At present, this scheduling mechanism can be implemented as external logic interfaced with the switch datapath.
Behavior Conventional CBS $\pi^{2}$ CBS
Transmission eligibility condition Credits $\geq 0$ Credits $\geq M$ (threshold)
Rate of credit accumulation when events are queued Credits increase at a rate defined by idleslope; user-specified constant A counter controls idleslope; each packet entering the queue increments it by 1
Rate of credit decrease during transmission Credits decrease at a rate defined by sendslope; user-specified constant Transmit one event and credits drop to 0. The sendslope parameter can be set very high to emulate this behavior
Credit reset behavior If the queue is empty, credits reset to 0 If the queue is empty, credits reset to 0
Table S1: Comparison of conventional CBS and $\pi^{2}$ CBS mechanisms.
Figure S4: A) Conventional CBS implementation architecture extracted from [cbs_google]. B) The $\pi^{2}$-CBS-based implementation—The modifications from the conventional implementation are highlighted in red. The implementation of a differential $\pi^{2}$ neuron (Eq. 6 of the manuscript) is presented here. As depicted, $\pi^{2}$ computation requires modest microarchitectural extensions to the CBS implementation—specifically programmable/non-zero credit thresholds (an extension discussed in [cbs_threshold]) and queue length-dependent credit accumulation—which extends the internal credit-update logic to support $\pi^{2}$ computation. These changes preserve IEEE standards and Ethernet frame formats.
1:/* Parameters set during configuration */
[…]
3:burstSize = 0
4:MaxTime = w
5:/* Initialization */
[…]
10:/* Frame Processing */
[…]
16:   $T_{\text{groupEligibility}},$
17:   $T_{\text{shaperEligibility}})$
18:/* Shaping */
19:if $T_{\text{eligibility}}\leq(T_{\text{arrival}}+MaxTime)$ then
20:  $T_{\text{groupEligibility}}=T_{\text{shaperEligibility}}$
21:  $T_{\text{bucketEmpty}}=(T_{\text{eligibility}}<T_{\text{bucketFull}})?$
22:     $T_{\text{shaperEligibility}}:$
23:     $T_{\text{shaperEligibility}}+T_{\text{eligibility}}-T_{\text{bucketFull}};$
24:  $AssignAndProcess(frame,T_{\text{arrival}}+MaxTime)$ (modified)
25:else
26:  $Discard(frame);$
27:end if
Algorithm 1 : In $\pi^{2}$-ATS, the standard ATS eligibility computation and residence-time admission check are preserved without modification. By configuring the committed rate to a large value and setting the burst size to zero, eligibility reduces to the arrival timestamp. After successful admission, frames are scheduled at the precomputed value (Tarrival+MaxTime), reinterpreting the existing ATS residence-time bound as a synaptic delay. As a result, there is no additional energy consumption beyond that of the conventional ATS datapath. No frame formats or scheduling rules defined by IEEE TSN are altered. While conventional ATS implementations transmit frames immediately upon eligibility, $\pi^{2}$-ATS deliberately schedules transmission at the residence-time bound (Tarrival+MaxTime), which remains fully compliant with IEEE TSN timing constraints. The conventional ATS protocol is adapted from [ats1, ats2], and the modifications required for $\pi^{2}$ processing (Eq.7,8) are highlighted with the keyword modified.
@@ 3.2 Vision for large-scale systems
Figure S5: A) A subpool of h neurons is fully connected to another subpool of h neurons. B) A port is shared by m neurons. C) The proposed $\pi^{2}$-NN implementation using an Ethernet switch. The Ethernet switch can support $n=mH_{out}$ neurons in total.
We strategize and analyze an implementation for a processing system consisting of n neurons, communicating through an Ethernet switch (with H ports each supporting a bandwidth B), as shown in Fig. S5A. $m$ neurons share a port. There are sub-pools (among these m neurons), each consisting of h neurons. Thus, a port is shared by m/h sub-pools (Fig. S5B). We propose an architecture where a sub-pool of h neurons is fully connected to another sub-pool of h neurons (Fig. S5C). Each sub-pool-pool connection consists of $h^{2}$ synapses. For each of these synaptic connections, p bits representing the synaptic delays have to be stored. The number of sub-pools in the core is $n/h$. The neuron’s address consists of the port address it corresponds to ($\log_{2}(n/m)\ \text{ bits}$), the pool it belongs to ($\log_{2}(m/h)\ \text{ bits}$), distinct neuron address in that pool ($\log_{2}(h)\ \text{bits}$) - which amounts to a total of $\log_{2}(n)\ \text{ bits}$. Features of this system include:
@@
  • The Tx address is decoded, the routing table is accessed, and events with appropriate destination addresses (Rx address) are generated. Similar to the previous implementation, a PCP tag (p bits) is added to the events representing the synaptic delay value. The events are directed to their corresponding destination ports by the traffic classification and scheduling logic.
  • Every output port is associated with $m$ set of shaped queues (Fig. S5D), one for each neuron that shares that port. A set of shaped queues consists of $2^{p}$ queues, each one for each distinct PCP value. Events associated with the same PCP tags for a neuron are dropped in the same queue. So, the events will inherently be arranged in the ascending order of their arrival times. Thus, the queue selection is based on the Rx neuron address and the PCP tag.
  • The ATS shaper controls each shaped queue. Based on the synaptic delay (PCP tag), a TET tag is assigned to the events to schedule their transmission to the shared queue of the neuron by the ATS shaper. When the appropriate TET value of an event is reached, it is released to the shared queue of that Rx neuron.
  • A modified CBS shaper controls the shared queue of the neuron. As events arrive, the credits are accumulated. When the CBS-controlled gate opens for transmission, an event encoded with the Rx neuron address, consisting of $log_{2}(n)$ bits, is forwarded to the output ports of the switch. The remaining events are dropped (TTFS encoding). This event now becomes the next Tx event processed by the switch.
@@
  • For the power analysis, we assume the weights are uniformly distributed (a similar assumption was taken in [truenorth]).
Bandwidth requirements at the input/output port: On average, the number of neurons (m) that can share a port or channel that can support a bandwidth of B bits/s can be computed by solving the following inequality.
$B\geq F_{spk}mlog_{2}(n)$ (S15)
Memory requirements: Say we have a look-up table in the switch storing the connectivity information of every pool-pool pair and their corresponding synaptic delays. Here, we assume that every input port is associated with a memory block storing the connectivity information of the m neurons that share that port. Once a Tx event arrives at an input port, its pool number is extracted ($log_{2}(m/h)\ bits$), and the LUT is accessed to retrieve the destination port ($log_{2}(n/m)\ bits$) and the destination pool address ($log_{2}(m/h)\ bits$) it is associated with. The Tx event is fully connected to every neuron in the destination pool. For every pool-pool connection, we have to store $h^{2}p$ bits. Thus, for one pool-pool connection, the LUT stores $h^{2}p+log_{2}(nm/h^{2})\ bits$. Every port is associated with m/h pools, and there are $H_{in}$ ports. Thus, the total memory required to store the LUT for all the input ports will be:
$M\geq nhp+(n/h)log_{2}(mn/h^{2})\ bits$ (S16)
@@
$Q_{nsize}=nKlog_{2}(n)F_{spk}T_{out}\ bits(time\ units/s)$ (S17)
@@
$QN_{size}=T_{avg}KF_{spk}(log_{2}(n)+n_{t})\ bits(time\ units/s)$ (S19)
Here, $n_{t}$ is the number of bits required to store the TET (transmission eligibility time). If more than $KF_{spk}\ events/s$ occur, then they will be automatically dropped. If the synaptic delay values are uniformly distributed among the $2^{p}$ queues of a neuron, then the queue size can be further reduced by $2^{p}$ times. Under this assumption, the total queue memory required in bytes across all the output ports is:
$Q_{size}=T_{avg}nKF_{spk}(log_{2}(n)+n_{t})/(8*2^{p})\ B(time\ units/s)$ (S20)
Throughput: Neuromorphic systems are usually compared in terms of synaptic operations per second (SOPS). The SOPS for this system is related to its network activity and can be computed as a product of the average firing rate of the neurons and the average number of active synapses [truenorth2, spinnaker]. The throughput of the $\pi^{2}$ system can be computed based on its three fundamental operations: number of effective delay operations performed, number of packets transmitted by the interconnect system (neuron activity), and number of packet drops at the queues of the output ports (at the neuron level).
@@
It is important to note that these operations are primarily implemented using the interconnect system (Ethernet switches, queues).
@@
Power analysis: The switches, interconnect transmission power, and memory access dominate the system’s network power consumption. If the interconnect modules to the switch consume aJ/bita\ J/bit then the interconnect power can be calculated as:
$P_{i}=aBH\ Watts$ (S25)
@@
If it takes bJ/bitb\ J/bit to access the memory related to the routing table and queues at the output port, then the power consumption to access Vbits/sV\ bits/s is
$P_{m}=bV\ Watts$ (S26)
@@
The total power of the system, consisting of a switch, memory, and interconnects (excluding the neurons), can be computed as:
$P_{total}=P_{s}+P_{i}+P_{m}\ Watts$ (S27)